OpenStudy (anonymous):

integrate 1/((x-3)^2 +1)

OpenStudy (anonymous):

i sorta know how to do it but I'm not sure why theres a negative in front of the tan^-1

OpenStudy (anonymous):

\[\int\frac{1}{(x-3)^2+1} dx\] ?

OpenStudy (anonymous):

yeah thats it

OpenStudy (anonymous):

u sub is the way to go u= x-3 du= dx \[\int\frac{1}{u^2+1} du\]

OpenStudy (anonymous):

yep got that far but I'm missing a negative sign infront of my answer lol

OpenStudy (anonymous):

we know this is equal to \[arctan(u)+C\]

OpenStudy (anonymous):

u=x-3 hence our answer is \[arctan(x-3)+C\]

OpenStudy (anonymous):

there is a negative sign in front of it in the answers

OpenStudy (anonymous):

\[-arctan(3-x)=arctan(x-3)\]

OpenStudy (anonymous):

oh yeah lol thats right my bad i ended up with the right answer. thanks:)